OpenStudy (anonymous):

solve for x: 3/3x+1/x+4=10/7x

OpenStudy (anonymous):

x=-1/7

OpenStudy (anonymous):

2/x + 4 = 10/7x 2*7+4*7x=10 28x=-4 x=-1/7